# Convergence for Natural Policy Gradient on Infinite-State Queueing MDPs

### v2

@@ 1. Introduction
This paper focuses on average-reward performance objectives for queueing problems, rather than discounted-reward objectives, as average-reward objectives are most often of importance to queueing practitioners. While infinite-state discounted-reward MDPs have rarely been explicitly studied, results on the NPG algorithm in such settings follow from existing finite-state NPG results in the discounted setting, given the special structure of queueing MDPs. This is in contrast to infinite-state average-reward results, for which new techniques were needed. We discuss infinite-state discounted-reward MDPs in Section 6.
@@
  • Section 6: We briefly discuss the infinite-state discounted-reward setting.
@@ 2.1. Natural Policy Gradient with finite state-space
As we discuss in Section 6, this result straightforwardly generalizes from the finite-state-space setting with discounted reward to the infinite-state-space setting with discounted reward, given mild assumptions on the structure of the infinite-state MDP, proving the same $O(1/T)$ convergence result.
@@ 2.5. Concurrent Work
In concurrent work currently under submission, Murthy et al. (Murthy et al., 2024) study a distinct, but related setting using the same framework as in this paper. In this paper, we study queueing MDPs in a noise-free setting, where the NPG updates can be computed exactly. Murthy et al. (Murthy et al., 2024) maintain the same queueing MDPs and generalize the setting to allow noisy NPG updates, but examine a restricted policy space, where all policies in the space are “uniformly stable”: All policies satisfy the Foster-Lyapunov theorem (Foster, 1953) with the same test function $V(s)$, the same bounded set, and the same drift $\epsilon$. Neither result is a strengthening of the other: This paper examines a general policy space under exact NPG updates, while the Murthy et al. (Murthy et al., 2024) paper allows noisy NPG updates but requires a restricted policy space with uniform stability.
@@ 3.1. Markov Decision Processes
We consider Markov Decision Processes (MDPs) with infinite horizon, infinite state space $S$, and finite action space $A$, and some class of randomized policies $\pi\in\Pi$. Each policy $\pi\in\Pi$ is a function from states to distributions over actions.
@@
$\mathbb{P}_{\pi}\{s^{\prime}\mid s\}:=\mathbb{P}\{s^{\prime}\mid s,\pi\}:=\sum\nolimits_{a\in A}\pi(a\mid s)\mathbb{P}\{s^{\prime}\mid s,a\}.$
The reward is given by a reward function $r(s,a)$ which maps a state-action pair to a real-valued reward. We likewise write $r(s,\pi)$ to denote the expected single-step reward associated with policy $\pi$ and state $s$. We define $r_{\max}(s)$ to be the maximum reward achievable in state $s$, $r_{\max}(s):=\max_{a}r(s,a)$. Note that we focus on reward-maximization rather than cost-minimization. Our primary reward function is the negation of the total queue length.
@@ 3.2. NPG Algorithm
Note that the NPG algorithm is an MDP optimization algorithm, rather than a learning algorithm. In particular, we assume that the relative value function $Q_{\pi}(s,a)$ can be exactly computed.
@@ 3.3. Definition of Generalized Switch with Static Environment
  • For each pair $(i,j)$, it is possible for more jobs arrive than depart ($P(w_{i}^{j}>v_{i})>0$).
  • For each pair $(i,j)$, it is possible for equally many jobs to arrive as depart ($P(w_{i}^{j}=v_{i})>0$).
@@
Our primary reward function $r(s,\cdot)$ of state $s=\{q_{i}\}$ under any action is the negative total queue length $r(s,\cdot)=-\sum_{i}q_{i}=-|q|_{1}$. We will also consider the setting where the reward is the negative $\alpha$-th moment of the total queue length, $-(\sum_{i}q_{i})^{\alpha}$, for a generic $\alpha\geq 1$.
@@ 3.4. MaxWeight
(3) $\text{MaxWeight}(q):=\arg\max_{j}\sum\nolimits_{i}q_{i}\mu^{j}_{i}$
@@ 4.1. Challenges
With respect to Even-Dar et al. (Even-Dar et al., 2009), combining that paper’s Algorithm 4 (MDP Experts) with its Algorithm 1 (Weighted Majority) results in exactly the Natural Policy Gradient algorithm, and the paper’s Theorem 4.1 proves that NPG converges to the optimal policy in the finite-state average-reward setting, with convergence rate $O(1/\sqrt{T})$. To prove this theorem, the paper assumes that there exists some finite mixing time $\tau$ for all policies $\pi$. Intuitively, this assumption states that for every policy and every initial system state, the system evolves to a state distribution becomes near the policy’s stationary distribution in at most $O(\tau)$ steps. In a finite-state MDP, this is a reasonable assumption: For reasonable policies, $\tau$ might be exponential in $|S|$ at worst. In contrast, in a queueing system with unbounded queue length, this assumption is plainly false. No policy can achieve a finite mixing time, much less all policies.
@@ 5.1. Background lemmas
(Cesa-Bianchi et al., 1997, Theorem 4.4.3) Consider an instance of the expert advice problem, where for any step $k$ and any pair of actions $a,a^{\prime}$, $M$ is an upper bound on the value of $\widetilde{r}_{k}(a)-r_{k}(a^{\prime})$. By selecting
@@ 5.2. Step 0: Stability
We will use the function $-V_{k}(s)$ as a Lyapunov function for the policy $\pi_{k+1}$, and apply the Foster-Lyapunov theorem (Foster, 1953). To show that $\pi_{k+1}$ is a stable policy with $-V_{k}(s)$ as our Lyapunov function, we must show that
  • $V_{k}(s)$ is upper bounded by a constant for all $s$,
  • There exists a constant $C$ such that for all $s$, the expected one-step update of $V_{k}(s)$ is lower bounded:
    $\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)\geq C,$
  • There exists a finite set $S$ and positive constant $\epsilon>0$ such that the expected one-step update of $V_{k}(s)$ is bounded above zero:
    $\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)\geq\epsilon.$
Let’s start by upper bounding $V_{k}(s)$. Note that the maximum value of $V_{k}(s)$ must be attained by a state $s$ such that the maximum possible reward $r_{\max}(s)$ exceeds $J_{k}$, as can be seen by examining the Poisson equation (1). There are only finitely many such states, by Assumption 1(b), so $V_{k}(s)$ is upper bounded.
@@
$\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,a)\geq J_{k}-c_{\max}$
$\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,a)\geq J_{k}-(J_{k}-1)=1.$
Thus, all conditions on $V_{k}(s)$ are satisfied, and by the Foster-Lyapunov theorem, policy $\pi_{k+1}$ is stable. By induction, all NPG iterates are stable. ∎
@@ 5.3. Step 1: Bounding change in relative value under NPG update
Knowing that the expected mean reward $J_{\pi}$ is at least $y$, we can apply renewal reward:
@@
$\tau_{\pi}(s)\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))({T^{s}_{out}}^{\prime}+\tau_{\pi}(s^{\prime}))$
where $s^{\prime}$ is the state at which we re-enter $\mathbb{S}_{z}$. Thus,
$\tau_{\pi}(s)$ $\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))\left(\frac{T^{s}_{out}}{1-p_{\pi,z}(s)}+\tau_{\pi}(s^{\prime})\right)$
$=x_{z}+t_{\pi,z}(s)+(1-p_{\pi,z}(s))(\tau_{\pi}(s^{\prime})-x_{z})$
@@ 5.5. Step 3: Bounding relative value $V_{k}(s)$ for all states $s$
Recall that by Lemma 3, $J_{\pi}\geq J_{0}>\pi$, so $-J_{k}+z<0$, so the direction of the inequality flips. Here we use Lemma 5, which states that all NPG iterates are stable.
The key fact relating $V_{k}(s)$ and $V_{k+1}(s)$ is Lemma 6: $V_{k+1}(s)\geq V_{k}(s)-\tau_{k+1}(s)(J_{k+1}-J_{k})$. Combining our bounds, we find that
@@ 5.6. Step 4: Proof of general result
(10) $\beta_{s}:=g\left(\sqrt{\frac{\ln|A|}{TM_{s}}}\right),\text{where}g(z)=1+2z+z^{2}/\ln 2.$
@@
$\sum\nolimits_{k=1}^{T}Q_{k}(s,\pi_{*})-V_{k}(s)$ $\leq\sqrt{TM_{s}\ln|A|}+\log_{2}(|A|)$
$=\sqrt{T\ln|A|}\sqrt{c_{2}|\hat{r}_{\max}(s)^{2}|+c_{3}|\hat{r}_{\max}(s)|+c_{4}}+\log_{2}(|A|)/2$
$\leq\sqrt{T\ln|A|}(c_{5}|\hat{r}_{\max}(s)|+c_{6})+\log_{2}(|A|)/2$
@@
$\sum\nolimits_{k=1}^{T}J_{*}-J_{k}$ $\leq\sqrt{T\ln|A|}(c_{5}E_{s\sim\pi_{*}}[c_{\max}-r(s,\pi_{*})]+c_{6})+\log_{2}(|A|)/2$
$=\sqrt{T\ln|A|}(c_{5}(c_{\max}-J_{*})+c_{6})+\log_{2}(|A|)/2$
@@
$\sum\nolimits_{k=1}^{T}J_{*}-J_{k}\leq c_{*}\sqrt{T},\text{ where}c_{*}:=\sqrt{\ln|A|}(c_{5}(c_{\max}-J_{*})+c_{6})+\log_{2}(|A|)/2.$

### v3

# Convergence of Natural Policy Gradient for a Family of Infinite-State Queueing MDPs

@@ 1. Introduction
This paper focuses on average-reward performance objectives for queueing problems, rather than discounted-reward objectives, as average-reward objectives are most often of importance to queueing practitioners. While infinite-state discounted-reward MDPs have rarely been explicitly studied, results on the NPG algorithm in such settings follow from existing finite-state NPG results in the discounted setting, given the special structure of queueing MDPs. This is in contrast to infinite-state average-reward results, for which new techniques were needed. We demonstrate the convergence of NPG for infinite-state discounted-reward MDPs in Section 6, generalizing a standard finite-state proof.
As a disclaimer, this paper does not consider the practical considerations of function approximation and computational complexity. This paper focuses on a setting where exact state-action relative values are available for every state, and studies the exact NPG algorithm, which manipulates and updates a policy with an infinitely long description. We leave the problems of analyzing the effects of approximate state-action relative value and analyzing computationally efficient variants of the NPG algorithm to concurrent and future work. See Section 2.5 for our concurrent work on function approximation in the infinite-state setting. One approach to achieving computational efficiency is to incorporate a neural-network-based policy approximation, which we leave to future work.
@@
  • Section 6: In the discounted-reward setting, we generalize a standard finite-state convergence proof to the infinite-state setting.
@@ 2.1. Natural Policy Gradient with finite state-space
As we discuss in Section 6, this result generalizes from the finite-state-space setting with discounted reward to the infinite-state-space setting with discounted reward, given mild assumptions on the structure of the infinite-state MDP, proving the same $O(1/T)$ convergence result.
@@ 2.5. Concurrent Work
In concurrent work currently under submission, Murthy et al. (Murthy et al., 2024) study a distinct, but related setting using the same framework as in this paper.
That paper also considers a class of infinite-state-space finite-action-space MDPs, with the objective of maximizing average reward. The class of MDPs considered in that work is more general in structure than the MDPs considered in this paper, but the policy space is more restricted. That paper does not make the structural assumption Assumption 1 and uses a modified version of the connectivity assumption Assumption 2. Instead, that paper makes three assumptions:
  1. (1)
    For all policies $\pi$ in the policy space $\Pi$, the induced Markov chain is irreducible.
  2. (2)
    There exists a fixed Lyapunov function $f$ and a fixed finite set $B$ such that for every policy $\pi\in\Pi$, the function $f$ and finite set $B$ prove the stability of the underlying Markov chain via the Foster-Lyapunov theorem (Foster, 1953; Brémaud, 1999), in a uniform fashion.
  3. (3)
    The finite set $B$ is uniformly connected under all policies $\pi$, similarly to Assumption 2.
Note that Item 2 in the concurrent paper implies that all policies $\pi\in\Pi$ are stable, and further implies a uniform lower bound on the average reward of all policies $\pi\in\Pi$. This paper makes no such assumption.
Furthermore, the concurrent paper studies the noisy NPG algorithm, where function approximation is used to estimate the state-action relative value $Q_{k}(s,a)$, and policy updates are performed using the approximate relative value. In contrast, our paper only studies the noisy-free NPG algorithm. Both papers set the step-size $\beta_{s}$ to depend on the state $s$.
Neither result is a strengthening of the other: This paper examines a general policy space under noise-free NPG updates, while the Murthy et al. (Murthy et al., 2024) paper allows noisy NPG updates but requires a restricted policy space with uniform stability.
@@ 3.1. Markov Decision Processes
We consider Markov Decision Processes (MDPs) with infinite horizon and infinite state space $S$. We assume that for each state $s\in S$, there exists a finite action space $A_{s}$, with a uniform bound on the size of the action space $|A_{s}|\leq A_{\max}$. We define the policy space $\Pi$ to be the set of all functions $\pi$ from states to distributions over actions.
@@
$\mathbb{P}_{\pi}\{s^{\prime}\mid s\}:=\mathbb{P}\{s^{\prime}\mid s,\pi\}:=\sum\nolimits_{a\in A_{s}}\pi(a\mid s)\mathbb{P}\{s^{\prime}\mid s,a\}.$
The reward is given by a reward function $r(s,a)$ which maps a state-action pair to a real-valued reward. We likewise write $r(s,\pi)$ to denote the expected single-step reward associated with policy $\pi$ and state $s$:
$r(s,\pi)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)]$
We define $r_{\max}(s)$ to be the maximum reward achievable in state $s$, $r_{\max}(s):=\max_{a}r(s,a)$. Note that we focus on reward-maximization rather than cost-minimization. Our primary reward function is the negation of the total queue length.
@@ 3.2. NPG Algorithm
Note that the NPG algorithm is an MDP optimization algorithm, rather than a learning algorithm. In particular, we assume that the relative value function $Q_{\pi}(s,a)$ can be exactly computed, rather than approximated, and we do not consider the practical implementation of the algorithm, leaving the problem of making the algorithm computationally efficient to future work.
@@ 3.3. Definition of Generalized Switch with Static Environment
Our primary reward function $r(s,\cdot)$ of state $s=\{q_{i}\}$ under any action is the negative total queue length $r(s,\cdot)=-\sum_{i}q_{i}=-|q|_{1}$. We will also consider the setting where the reward is the negative $\alpha$-th moment of the total queue length, $-(\sum_{i}q_{i})^{\alpha}$, for a generic $\alpha\geq 1$. Note that the specially-identified maximum reward state $\vec{0}$ is the empty state, where $\forall i,q_{i}=0$.
@@
  • For each pair $(i,j)$, it is possible for more jobs arrive than depart ($P(v_{i}>w_{i}^{j})>0$).
  • For each pair $(i,j)$, it is possible for equally many jobs to arrive as depart ($P(v_{i}=w_{i}^{j})>0$).
@@
More specifically, we define our action space $A_{s}$ in a given state $s=\{q_{i}\}$ to be the set of non-idling service options $j$ for the given state $s$, for any state $s$ which is not the empty state:
$A_{\{q_{i}\}}:=\{j\mid\exists i:q_{i}>0\land P(w_{i}^{j}>0)>0\},\forall\{q_{i}\}\neq\vec{0}$
As a special case, if $\{q_{i}\}$ is the empty state $\vec{0}$, any service option may be used: idling is unavoidable. The uniform bound on the size of the action space is satisfied with $A_{\max}=m$.
@@ 3.4. MaxWeight
(3) $\text{MaxWeight}(\{q_{i}\}):=\arg\max_{j}\sum\nolimits_{i}q_{i}\mu^{j}_{i}$
Note that MaxWeight always selects a non-idling service option $j$ for any state $\{q_{i}\}\neq\vec{0}$. This service option $j$ thus lies within our action space $A_{\{q_{i}\}}$.
@@ 3.5. Assumptions
Note that for Assumption 2 to hold, Assumption 1(b) is a necessary prerequisite assumption. Uniform connectedness of high-reward states is only possible if there are finitely many high-reward states.
@@ 4. Results
Our quadratic-growth-rate assumption in (5) on the initial policy is the most general assumption that suffices for our proof technique. Intuitively, a key step in the proof of Theorem 2 depends on proving that the expected square-root of the relative value of each NPG iterate $V_{k}(s)$ is finite, where the state $s$ is distributed according to the stationary distribution of the optimal policy $\pi^{*}$: $E_{s\sim\pi^{*}}[\sqrt{V_{k}(s)}]<\infty$. Using the quadratic-growth-rate assumption in (5), we can prove that $V_{k}(s)$ also grows at most quadratically for each NPG iterate, and thereby bound the expectation in comparison to the mean reward of the optimal policy, which is known to be small. Without (5), or with a weaker assumption, we would have no way to prove that this expectation is finite. See (11) and the accompanying discussion in the proof of Theorem 2 for more detail.
@@ 4.1. Challenges
With respect to Even-Dar et al. (Even-Dar et al., 2009), combining that paper’s Algorithm 4 (MDP Experts) with its Algorithm 1 (Weighted Majority) results in exactly the Natural Policy Gradient algorithm, and the paper’s Theorem 4.1 proves that NPG converges to the optimal policy in the finite-state average-reward setting, with convergence rate $O(1/\sqrt{T})$. To prove this theorem, the paper assumes that there exists some finite mixing time $\tau$ for all policies $\pi$. Intuitively, this assumption states that for every policy and every initial system state, the system evolves to a state distribution near the policy’s stationary distribution in at most $O(\tau)$ steps. In a finite-state MDP, this is a reasonable assumption: For reasonable policies, $\tau$ might be exponential in $|S|$ at worst. In contrast, in a queueing system with unbounded queue length and bounded completion rate, this assumption is plainly false. No policy can achieve a finite mixing time, much less all policies.
@@ 5.1. Background lemmas
(Cesa-Bianchi et al., 1997, Theorem 4.4.3) Consider an instance of the expert advice problem, where for any step $k$ and any pair of actions $a,a^{\prime}\in A$, $M$ is an upper bound on the value of $\widetilde{r}_{k}(a)-r_{k}(a^{\prime})$. By selecting
@@ 5.2. Step 0: Stability
We will use the function $b(s):=-V_{k}(s)$ as a Lyapunov function for the policy $\pi_{k+1}$, and apply the Foster-Lyapunov theorem (Foster, 1953; Brémaud, 1999). To show that $\pi_{k+1}$ is a stable policy with $b(s)$ as our Lyapunov function, we must show that
  • $b(s)$ is lower bounded by a constant for all $s$,
  • There exists a constant $C$ such that for all $s$, the expected one-step update of $b(s)$ is upper bounded:
    $\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[b(s^{\prime})]-b(s)\leq C,$
  • There exists a finite set $S$ and positive constant $\epsilon>0$ such that the expected one-step update of $b(s)$ is bounded below zero:
    $\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[b(s^{\prime})]-b(s)\leq-\epsilon.$
Translating these requirements on $b(s)=-V_{k}(s)$ into requirements on $V_{k}(s)$, we must:
  • upper bound $V_{k}(s)$ over all states $s$,
  • lower bound the expected one state update of $V_{k}(s)$ over all states $s$, and
  • lower bound the expected one state update of $V_{k}(s)$ over all states $s\not\in S$ as above $\epsilon>0$.
Let’s start by lower bounding $b(s)$, or equivalently upper bounding $V_{k}(s)$. Note that the maximum value of $V_{k}(s)$ must be attained by a state $s$ such that the maximum possible reward $r_{\max}(s)$ exceeds $J_{k}$, as can be seen by examining the Poisson equation (1). There are only finitely many such states, by Assumption 1(b), so $V_{k}(s)$ is upper bounded. Note that $J_{k}$ is finite because $\pi_{k}$ is stable by our inductive assumption.
@@
$\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,\pi_{k+1})\geq J_{k}-c_{\max}$
$\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,\pi_{k+1})\geq J_{k}-(J_{k}-1)=1.$
By lower bounding the expected change of $V_{k}(s)$, we equivalently upper bound the expected change of $b(s)=-V_{k}(s)$:
$\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[b(s^{\prime})]-b(s)$ $\leq c_{\max}-J_{k}$
$\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[b(s^{\prime})]-b(s)$ $\leq-1.$
Setting $C=c_{\max}-J_{k}$ and $\epsilon=1$, all conditions on $b(s)$ are satisfied, and by the Foster-Lyapunov theorem, policy $\pi_{k+1}$ is stable. By induction, all NPG iterates are stable. ∎
@@ 5.3. Step 1: Bounding change in relative value under NPG update
Knowing that the expected mean reward $J_{\pi}$ is at least $y$, we apply the Renewal-Reward theorem (Harchol-Balter, 2013, Theorem 23.4):
@@
$\tau_{\pi}(s)\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))({T^{s}_{out}}^{\prime}+\mathbb{E}[\tau_{\pi}(s^{\prime})]),$
where $s^{\prime}$ is a random variable denoting the state at which we re-enter $\mathbb{S}_{z}$. Thus,
$\tau_{\pi}(s)$ $\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))\left(\frac{T^{s}_{out}}{1-p_{\pi,z}(s)}+\mathbb{E}[\tau_{\pi}(s^{\prime})]\right)$
$=p_{\pi,z}(s)x_{z}+T^{s}_{out}+(1-p_{\pi,z}(s))\mathbb{E}[\tau_{\pi}(s^{\prime})]$
$=x_{z}+T^{s}_{out}+(1-p_{\pi,z}(s))(\mathbb{E}[\tau_{\pi}(s^{\prime})]-x_{z})$
@@ 5.5. Step 3: Bounding relative value $V_{k}(s)$ for all states $s$
Recall that by Lemma 3, $J_{\pi}\geq J_{0}>z$, so $-J_{k}+z<0$, so the direction of the inequality flips. Here we use Lemma 5, which states that all NPG iterates are stable.
The key fact relating $V_{k}(s)$ and $V_{k+1}(s)$ is Lemma 6: $V_{k+1}(s)\geq V_{k}(s)-\tau_{k+1}(s)(J_{k+1}-J_{k})$.
Combining our bounds, we find that
@@ 5.6. Step 4: Proof of general result
(10) $\beta_{s}:=g\left(\sqrt{\frac{\ln|A_{s}|}{TM_{s}}}\right),\text{where }g(z)=1+2z+z^{2}/\ln 2.$
@@
$\sum\nolimits_{k=1}^{T}Q_{k}(s,\pi_{*})-V_{k}(s)$ $\leq\sqrt{TM_{s}\ln|A_{s}|}+\log_{2}(|A_{s}|)$
$=\sqrt{T\ln|A_{s}|}\sqrt{c_{2}|\hat{r}_{\max}(s)^{2}|+c_{3}|\hat{r}_{\max}(s)|+c_{4}}+\log_{2}(|A_{s}|)/2$
$\leq\sqrt{T\ln A_{\max}}(c_{5}|\hat{r}_{\max}(s)|+c_{6})+\log_{2}(A_{\max})/2$
@@
$\sum\nolimits_{k=1}^{T}J_{*}-J_{k}$ $\leq\sqrt{T\ln A_{\max}}(c_{5}E_{s\sim\pi_{*}}[c_{\max}-r(s,\pi_{*})]+c_{6})+\log_{2}(A_{\max})/2$
$=\sqrt{T\ln A_{\max}}(c_{5}(c_{\max}-J_{*})+c_{6})+\log_{2}(A_{\max})/2$
@@
$\sum\nolimits_{k=1}^{T}J_{*}-J_{k}\leq c_{*}\sqrt{T},\text{ where}c_{*}:=\sqrt{\ln A_{\max}}(c_{5}(c_{\max}-J_{*})+c_{6})+\log_{2}(A)/2.$